OpenStudy (anonymous):

If A and B are independent events, which equation must be true? AA.)P(B|A) = P(A) B.)P(B|A) = P(B) C.)P(B|A) = P(A) × P(B) D.)P(B|A) = P(A) + P(B)

OpenStudy (jango_in_dtown):

If the events A and B are independent then P(AB)=P(A)P(B) Now P(B|A)=P(AB)/P(A)=P(A)P(B)/P(A)=P(B)

OpenStudy (jango_in_dtown):

Hence 2 is satisfied @tiannataylor890